# ReCode: Unify Plan and Action for Universal Granularity Control

### v2

@@ A.1 Datasets
ALFWorld (Shridhar et al., 2021) provides a text-only game environment where the agent is required to complete a series of long-horizon household tasks based on natural language instructions. Built upon the embodied vision benchmark ALFRED (Shridhar et al., 2020), ALFWorld inherits its six core task types: pick and place, examine in light, clean and place, heat and place, cool and place, and pick two and place. After executing a command (e.g., “go to countertop 1”), the agent receives immediate textual feedback from the environment (e.g., “You arrive at countertop 1. On the countertop 1, you see an egg 3, a peppershaker 1, and a tomato 1”). ALFWorld only provides a binary reward (1/0) to indicate whether the task was successfully completed.
@@
WebShop Yao et al. (2022) is an online shopping website environment constructed with 1.18 million real-world product data points. The agent must navigate the site by searching and clicking buttons to browse various products, with the goal of finding and purchasing a specific item that meets the given requirements. We follow the data partitioning from MPO (Xiong et al., 2025), using a test set of 200 instances. WebShop provides a dense reward between 0 and 1 based on the agent’s actions and interactions with products. A task is considered successful only if the reward is strictly equal to 1.
@@
ScienceWorld is a text-based simulation environment centered on completing basic scientific experiments. The agent needs to interact with the environment to measure certain values or to complete an experiment by applying scientific knowledge. The tasks in ScienceWorld are meticulously categorized, with each task having several variations and offering numerous different sub-goals. We adopt the data split from ETO (Song et al., 2024) and exclude Task-9 and Task-10 from the trajectories. ScienceWorld provides a dense reward between 0 and 1 for each task, based on the completion of its sub-goals.
@@ A.3.1 WebShop Examples
A.4 Case Study

### v3

@@ A.1 Datasets
ALFWorld (shridhar2021alfworld) provides a text-only game environment where the agent is required to complete a series of long-horizon household tasks based on natural language instructions. Built upon the embodied vision benchmark ALFRED (shridhar2020alfred), ALFWorld inherits its six core task types: pick and place, examine in light, clean and place, heat and place, cool and place, and pick two and place. After executing a command (e.g., “go to countertop 1”), the agent receives immediate textual feedback from the environment (e.g., “You arrive at countertop 1. On the countertop 1, you see an egg 3, a peppershaker 1, and a tomato 1”). ALFWorld only provides a binary reward (1/0) to indicate whether the task was successfully completed.
@@
WebShop yao2022webshop is an online shopping website environment constructed with 1.18 million real-world product data points. The agent must navigate the site by searching and clicking buttons to browse various products, with the goal of finding and purchasing a specific item that meets the given requirements. We follow the data partitioning from MPO (xiong2025mpo), using a test set of 200 instances. WebShop provides a dense reward between 0 and 1 based on the agent’s actions and interactions with products. A task is considered successful only if the reward is strictly equal to 1.
@@
ScienceWorld is a text-based simulation environment centered on completing basic scientific experiments. The agent needs to interact with the environment to measure certain values or to complete an experiment by applying scientific knowledge. The tasks in ScienceWorld are meticulously categorized, with each task having several variations and offering numerous different sub-goals. We adopt the data split from ETO (song2024trial) and exclude Task-9 and Task-10 from the trajectories. ScienceWorld provides a dense reward between 0 and 1 for each task, based on the completion of its sub-goals.
@@
ManiSkill
We evaluate embodied decision making using two manipulation tasks from ManiSkill (tao2025maniskill), PickCube v1 and PullCube v1, both implemented with the official Franka Panda robot. Each environment exposes a continuous control interface based on incremental end–effector pose and gripper commands. To make this interface compatible with LLMs, we wrap the ManiSkill environments and provide a small set of textual control actions with continuous parameters. These actions allow the LLM to issue fine grained motor commands through executable code.
We expose four textual actions that map directly to the underlying ManiSkill action vector: move(dx, dy, dz) specifies continuous translation of the end effector, rotate(d_roll, d_pitch, d_yaw) specifies continuous rotation increments, and gripper(opening) sets a continuous gripper opening value in . All parameters are real-valued floats and are clipped to the action bounds enforced by ManiSkill. This interface preserves the continuous nature of the control problem. Besides, to provide a compact textual observation to the LLM, we extract and serialize key environment states including the end–effector position, object position, goal position, and (for PickCube) the grasp status. This produces a consistent text description for each step while avoiding direct exposure of raw simulator states.
@@ A.3 Fewshot Examples
A.4 Quantitative Analysis of ReCode
To provide quantitative evidence for how ReCode adjusts its decision granularity during execution, we report statistics computed over trajectories on the unseen splits from ALFWorld, ScienceWorld and WebShop using GPT-4o mini. These metrics characterize the hierarchical structure that emerges from ReCode’s recursive expansion process.
We measure three structural properties of the generated decision trees and these metrics capture the balance between coarse and fine decisions across tasks.
Average recursion depth, defined as the maximum expansion depth reached in an episode, averaged over all trajectories;
Average branching factor, defined as the average number of child functions or substeps produced by each expansion;
Placeholder ratio, defined as the fraction of generated functions that remain as high-level placeholders rather than being expanded into primitive actions.
| Environment | Avg. Depth () | Avg. Branching () | Placeholder Ratio |
| ALFWorld | 0.19 | ||
| WebShop | 0.17 | ||
| ScienceWorld | 0.11 |
Discussion
The statistics reveal clear differences in how ReCode structures its decisions across environments. ALFWorld has the smallest average depth and branching factor, suggesting that its tasks rarely require deeply nested plans and instead emphasize structured exploration and local interaction in relatively simple scenes. WebShop uses deeper trees with larger branching factors, which is consistent with multi step product search, filtering, and comparison that naturally induce several layers of decision points. ScienceWorld exhibits the deepest and most complex decision structures, reflecting its multi stage experimental procedures and frequent decomposition into intermediate subgoals. Across all environments, the placeholder ratios remain relatively low and tend to decrease as task complexity increases. A plausible explanation is that more complex tasks require richer environment dynamics to support reliable planning, so ReCode executes more primitive actions and interacts more frequently with the environment before or alongside high level abstractions. These patterns provide quantitative evidence that ReCode adapts its decision structure and effective granularity to the structure and complexity of each environment.
A.5 Case Study
Initially, the agent formulates a hybrid plan that integrates abstract sub-tasks with concrete, primitive actions. This initial plan, shown in the following code block, includes distinct sub-procedures like find_and_take for complex procedures that require further decomposition, while directly embedding known, simple steps like run('go to dresser 1') as primitive actions within the main strategic flow.
The agent then executes this hybrid plan sequentially. Primitive actions like run(...) are executed directly by the environment interpreter. In contrast, placeholder functions such as find_and_take trigger an on-demand expansion, where the agent generates the necessary code to complete that specific sub-task just in time. This hybrid strategy allows the agent to maintain a high-level, strategic view of the task while committing to concrete actions only when the steps are simple and certain. It combines the robustness of hierarchical planning with the efficiency of direct execution, providing a flexible and powerful problem-solving framework that can adapt its level of abstraction as needed.